OpenStudy (rileyhansen):

The ordered pairs (2 -21) and (5 -45) are solutions to which of the following equations

OpenStudy (zyberg):

We need to know which equations you can choose from ;)

OpenStudy (rileyhansen):

HaHa Silly me!

OpenStudy (rileyhansen):

y = 3x - 5 y = -3x + 5 y = 3x + 5 y = -3x - 5

OpenStudy (er.mohd.amir):

put value of x and y and which one satisfy is Ur ans

OpenStudy (rileyhansen):

Please talk like a human I cant understand giberish.

OpenStudy (zyberg):

Try putting the first element of your pair as x and the second element in your pair as y to all equations. The ones that satisfy both of the pairs will be your answer.

OpenStudy (er.mohd.amir):

first pair (2 , -21) that is x=2 and y=-21 now take first eq y=3x-5 put value -21=3*2-5 -21=6-5 -21=1 which is false so chech all other eq which is give true is ur ans

OpenStudy (rileyhansen):

Ok fellas, I dont mean to be like this but I'm kind of in a predicament so an awnser would be great! I understand how to do it but I just need the answer.

OpenStudy (rileyhansen):

Hello! Anyone please!

OpenStudy (er.mohd.amir):

no eq satisfy pair so either eq are wrong of ordered pair